# Model-Based Analysis of Mining Fairness in a Blockchain

### v2

@@ Abstract
Mining fairness in blockchain refers to equality between the computational resources invested in mining and the block rewards received. There exists a dilemma wherein increasing the transaction processing capacity of a blockchain compromises mining fairness, which consequently undermines its decentralization. This dilemma remains unresolved even with methods such as the greedy heaviest observed subtree (GHOST) protocol, indicating that mining fairness is an inherent bottleneck in the transaction processing capacity of the blockchain system. However, despite its significance, there have been insufficient research studies that have quantitatively analyzed mining fairness. In this paper, we propose a method for calculating mining fairness. First, we approximated a complex blockchain network using a simple mathematical model, assuming that no more than two blocks are generated per round. Within this model, we quantitatively determined local mining fairness and derived several measures of global mining fairness based on local mining fairness. Subsequently, we validated by blockchain network simulations that our calculation method computes mining fairness in networks much more accurately than existing methods. Finally, we analyzed various networks from the perspective of mining fairness.
@@ 1 Introduction
Blockchain is a foundational technology used primarily in decentralized currency systems such as Bitcoin [1]. In blockchain systems, transactions are processed in units known as blocks. The creation of a block involves numerous hash calculations, a process referred to as mining. Nodes that perform mining are called miners. Each miner follows a fork choice rule to identify and extend the main chain. When miners successfully create a block, they may be rewarded via a coinbase transaction, by which they would receive what is known as a block reward. However, these block rewards are obtainable only when the blocks that have been generated become part of the main chain.
Mining fairness refers to equality between the computational resources invested in mining and the resulting block rewards. In other words, it is equality between the proportion of hashrate and the proportion of block rewards (hereafter referred to as the block reward rate). If all blocks were incorporated into the main chain, mining fairness would be achieved because the number of blocks generated by each miner would not be affected by the state of the network. However, in practice, not all blocks are included in the main chain because of blockchain forks, and when blocks are discarded as a result, mining fairness is compromised. Forks can be classified into two types: intentional (malicious) and unintentional. The latter occurs when multiple blocks are generated almost simultaneously. This study addresses mining fairness in the context of unintentional forks.
@@
Mining fairness introduces a tradeoff between the transaction processing capacity and decentralization in blockchain systems (Fig. 1); increasing the transaction processing capacity of a system compromises decentralization. The transaction processing capacity depends on the number of transactions processed per block and the block generation interval. To increase this capacity, one might increase the block size and reduce the block generation interval. However, it is well known that increasing block sizes and reducing block generation intervals result in higher fork rates [2]. As previously observed, an increase in the fork rate undermines mining fairness. If mining fairness is reduced, then some miner groups achieve higher profit rates than others. Consequently, miners with lower profit rates end up leaving the system, whereas those with higher profit rates expand, leading to centralization and reduced decentralization.
@@
The dilemma between transaction processing capacity and decentralization in blockchain systems that arises from mining fairness has not yet been resolved. In this context, mining fairness is an inherent bottleneck in the transaction processing capacity. Here, we demonstrate that the dilemma caused by mining fairness is inherent, using the countermeasures adopted by Ethereum [3] (modified greedy heaviest observed subtree (GHOST) protocol) as an example. Increasing the transaction processing capacity of a blockchain leads to more forks, which, in turn, causes two main problems. First, there is an increased risk of attacks such as 51% attacks and selfish mining [4, 5]. Second, mining fairness is compromised. To address the first problem, Ethereum has introduced the GHOST protocol. On the other hand, to address the second problem, which is regarding the impact on mining fairness, Ethereum partially rewards blocks that cause forks but are not incorporated into the main chain (stale blocks). However, this approach has its own problematic aspects. From another perspective, this implies that even blocks that cause forks can receive block rewards, thereby reinforcing economic incentives for attacks such as 51% attacks and selfish mining. This situation shares the same structure as the nothing-at-stake problem. In fact, it is known that in Ethereum, the risk of attacks that compromise mining fairness, including selfish mining [4], is increased [6, 7, 8]. This indicates that the measures taken by Ethereum do not fundamentally address the problems related to mining fairness.
@@
Thus, it is critical to perform further analyses on mining fairness. One possible approach to these analyses is to perform simulations, which, unfortunately, can be time-consuming and impractical. Consequently, alternative approaches have been explored [9, 10, 11, 12, 13, 14, 15, 16]. However, these methods do not correctly account for how mining fairness is compromised by forks (Section 6.1) and, consequently, lead to analyses that are based on only a weak reflection of real-world systems, making it difficult to derive meaningful solutions from them.
@@
In this study, we propose a model-based calculation method for quantitatively analyzing mining fairness. We approximate a blockchain network using a simplified model, assuming that each round contains at most two blocks. A round $r$ is defined as a unit of time that starts with the creation of a block at height $r$. Essentially, we also assume that at most one fork occurs per round. This modeling approach simplifies the analysis of complex blockchain networks.
@@
Afterward, we validate the accuracy of our proposed model-based calculation in measuring mining fairness by conducting simulation experiments. Owing to computational constraints, validating our method in large-scale networks would be challenging. Instead, we perform the validation in networks with small numbers of miners. Our results demonstrate that the model-based calculation quantitatively determines mining fairness much more accurately than existing methods.
The model-based calculation enables efficient and comprehensive analysis of mining fairness. In this study, we analyze mining fairness using the model-based calculation. First, we examine mining fairness in a simple network composed of two miners. Subsequently, we investigate the impact of hashrate distribution on mining fairness. Finally, we reassess existing blockchain network improvement methods from the perspective of mining fairness.
@@ 2 Rounds
Herein, to accurately capture and incorporate the impact of forks, we introduce the concept of rounds in a blockchain. A round is a time interval defined from a global perspective. Specifically, round $r$ refers to the time from the first generation of a block at height $r$ to the first generation of a block at height $r+1$. The height of round $r$ is defined as $r$.
@@
Because forks occur probabilistically within a blockchain network, the number of blocks per round is not always equal to one. After the first block at height $r$ has been generated, another miner may generate a new block before the first block is propagated throughout the network. It should be noted that not all blocks generated in round $r$ have a height of $r$. For example, a miner unaware of the block at height $r-1$ will generate a block at height $r-1$ during round $r$ (Fig. 2).
@@
The round start rate of each miner is defined as the probability that the miner initiates a round. Note that the round start rate does not equal the proportion of the hashrate, owing to the occurrence of forks.
@@
With the introduction of rounds, we can formally define the fork rate. The fork rate is the probability that the number of blocks per round is two or more. When the number of blocks per round is two or more, the blockchain diverges, or forks. Therefore, forks (and fork rates) are named as such because they function similarly to a traditional fork.
@@ 3.2 Definition of Mining Fairness
Before calculating mining fairness, we first have to properly define it. In this study, mining fairness is divided into local mining fairness and global mining fairness. We define local mining fairness, $LF$, based on two measures as follows:
@@ 3.3 Calculation of Mining Fairness
Let $X_{r}$ be a random variable representing the miner that generates the block that starts round $r$. Then, the following equation holds:
$P(X_{r+1}=i)=\sum_{j\in V}\left(\alpha_{i}(1-F_{ji})+\sum_{k}\alpha_{k}F_{jk}\alpha_{i}\right)P(X_{r}=j).$ (5)
Note that $P(X_{r+1}=i)$ is dependent only on $P(X_{r}=j)$. Therefore, the stochastic process $\{X_{r}\}_{r=0}^{\infty}$ is a Markov chain. Additionally, because $F_{ij}$ is less than 1 and $\alpha_{i}(1-F_{ji})+\sum_{k}\alpha_{k}F_{jk}\alpha_{i}$ is usually positive, this Markov chain is ergodic in most cases. Consequently, a unique stationary distribution exists, and the limiting distribution is stationary. We can then determine the stationary distribution by iterating (5).
Let the limiting distribution be $\pi$. This represents the distribution of the miners that generate blocks that start rounds after sufficient time has passed. Using $\pi$, the block reward rate $r_{i}$ for each miner is given by the following equation:
@@ 3.4 Algorithm
In this section, we describe the algorithm used in this study to calculate mining fairness, as detailed in Section 3.3.
Algorithm 1 employs an iterative method to compute the mining fairness of each miner. The round start rate calculation is performed calculated between lines 9 and 27. Specifically, The fork rate is precomputed between lines 10 and 16. The variable loop manages the operations executed in each iteration. The calculations within the for loop from lines 20 to 26 follow the same process as described in (5). Mining fairness is computed between lines 28 and 34, with the for loop calculations corresponding to equations (8) and (9).
@@
2:$V$triangleright set of miners
3:$N$triangleright number of miners
4:$\alpha[N]$triangleright proportion of hashrate
5:$F[N][N]$triangleright fork rate
6:$W[N][N]$triangleright winning rate
@@
9:$\pi[N][2]$triangleright round start rate
10:$LF_{1}[N]$triangleright $LF_{1}$
11:$LF_{2}[N]$triangleright $LF_{2}$
@@
14:$\epsilon$triangleright error
15:$dp[N]$triangleright for dynamic programming
@@
[…]
@@
[…]
@@
38:         $LF_{1}[i]\leftarrow LF_{1}[i]+\pi(j)\alpha_{i}F_{ji}(1-W_{ji})-\pi(i)\alpha_{j}F_{ij}(1-W_{ij})$
@@ 3.5.1 How to Determine $F_{ij}$
3.5.2 How to determine $W_{ij}$
Prior to any explanations regarding $W_{ij}$, it is first essential to discuss the concept of chain ties.
Each miner constructs chains from their blocks and selects one main chain among them. The rule for selecting this chain is known as the fork choice rule. In Bitcoin, for example, the longest chain rule, which selects the longest chain, is adopted.
However, in some cases, the fork choice rule alone may not uniquely determine the main chain owing to the occurrence of forks. This situation is known as a chain tie. To resolve a chain tie, a tie-breaking rule is implemented. In Bitcoin, for example, the first-seen rule, which selects the chain received first, is adopted.
@@ 3.5.2 How to determine $W_{ij}$
Selects the latest chain among the chains in a tie [17, 18]. Suppresses selfish mining more effectively than the random rule.
@@ 4 Validation
Herein, we validate the capability of our proposed model-based calculation to accurately determine mining fairness. First, we examine the assumption that the number of blocks per round is at most two from the perspective of the scale of the fork. Next, we compare the results of simulation experiments with those of the model-based calculation.
Whereas it is excessively time-consuming to calculate mining fairness via simulations of networks composed of many (approximately 100 or more) miners, it is feasible to calculate mining fairness accurately and relatively quickly for networks with a small number of miners (2–10). In this study, we perform the validation using networks consisting of two and ten miners and show that our proposed model-based calculation determines mining fairness much more accurately than existing methods.

### v3

# Model-Based Calculation Method of Mining Fairness in Blockchain

@@ Abstract
Mining fairness in blockchain refers to equality between the computational resources invested in mining and the block rewards received. There exists a dilemma wherein increasing the transaction processing capacity of a blockchain compromises mining fairness, thereby undermining its decentralization. This dilemma remains unresolved despite methods such as the greedy heaviest observed subtree (GHOST) protocol, indicating that mining fairness is an inherent bottleneck in the transaction processing capacity of the blockchain system. However, despite its significance, existing analyses neglect the impact of blockchain forks, resulting in imprecise evaluations and limited insights. To address this issue, we propose a method for calculating mining fairness that explicitly captures the influence of forks. First, we approximate a complex blockchain network using a simple mathematical model, assuming that no more than two blocks are generated per round. Within this model, we quantitatively determine local mining fairness and derive several measures of global mining fairness based on local mining fairness. Subsequently, we validated by blockchain network simulations that our calculation method computes mining fairness in networks much more accurately than existing methods. The proposed method facilitates a rigorous evaluation of trade-offs between scalability and decentralization by offering a clear, quantitative framework for measuring and comparing reward distribution among miners. Consequently, it is expected to provide valuable insights for future mining fairness research and the design of next-generation blockchain systems.
@@ 1 Introduction
Blockchain is a foundational technology primarily used in decentralized currency systems such as Bitcoin [1]. In blockchain systems, transactions are processed in units known as blocks. Generating a block involves numerous hash calculations, a process referred to as mining. Nodes that perform mining are called miners. Each miner follows a fork choice rule to identify and extend the main chain. When miners successfully generate a block, they may be rewarded via a coinbase transaction, by which they would receive what is known as a block reward. However, these block rewards are obtainable only when the blocks that have been generated become part of the main chain.
Mining fairness refers to equality between the computational resources invested in mining and the resulting block rewards; that is, it is equality between the proportion of hashrate and the proportion of block rewards (hereafter referred to as the block reward rate). If all blocks were incorporated into the main chain, mining fairness would be achieved because the number of blocks generated by each miner would not be affected by the state of the network. However, in practice, not all blocks are included in the main chain because of blockchain forks, and mining fairness is compromised when blocks are discarded. Forks can be classified into two types: intentional (malicious) and unintentional. The latter occurs when multiple blocks are generated almost simultaneously. This study addresses mining fairness in the context of unintentional forks.
@@
Mining fairness introduces a trade-off between the transaction processing capacity and decentralization in blockchain systems (Fig. 1)—increasing the transaction processing capacity of a system compromises decentralization. The transaction processing capacity depends on the number of transactions processed per block and the block generation interval. To increase this capacity, one might increase the block size and reduce the block generation interval. However, it is well known that increasing block sizes and reducing block generation intervals result in higher fork rates [2]. As observed previously, an increase in the fork rate undermines mining fairness. If mining fairness is reduced, some miner groups achieve higher profit rates than others. Consequently, miners with lower profit rates end up leaving the system, whereas those with higher profit rates expand, leading to centralization and reduced decentralization.
@@
The dilemma between transaction processing capacity and decentralization in blockchain systems that arises from mining fairness has yet to be resolved. In this context, mining fairness is an inherent bottleneck in the transaction processing capacity. Here, we demonstrate that the dilemma caused by mining fairness is inherent, using the countermeasures adopted by Ethereum [3] (modified greedy heaviest observed subtree (GHOST) protocol [4]) as an example. Increasing the transaction processing capacity of a blockchain leads to more forks, which, in turn, causes two main problems. First, there is an increased risk of attacks such as double-spending attacks and selfish mining [5, 6]. Second, mining fairness is compromised. To address the first problem, Ethereum has introduced the GHOST protocol. In addition, to address the second problem regarding the impact on mining fairness, Ethereum partially rewards blocks that cause forks but are not incorporated into the main chain (stale blocks). However, this approach has its own challenges. From another perspective, this implies that even blocks that cause forks can receive block rewards, thereby reinforcing economic incentives for attacks such as double-spending attacks and selfish mining. This situation shares the same structure as the nothing-at-stake problem. It is well known that, in Ethereum, the risk of attacks that compromise mining fairness, including selfish mining [5], is increased [7, 8, 9]; this indicates that the measures taken by Ethereum do not fundamentally solve the problems related to mining fairness.
@@
Thus, it is crucial to perform further analyses on mining fairness. One possible approach to these analyses is to perform simulations, which, unfortunately, is time-consuming and impractical. Consequently, alternative approaches have been explored [10, 11, 12, 13, 14, 15, 16, 17]. However, these methods do not accurately account for how mining fairness is compromised by forks (Section 3) and, consequently, lead to analyses based on only a weak reflection of real-world systems, making it challenging to derive meaningful insights from them.
@@
In this study, we propose a model-based calculation method for quantitatively analyzing mining fairness. We approximate a blockchain network using a simplified model, assuming that each round contains at most two blocks. A round $r$ is defined as a unit of time that starts with the generation of a block at height $r$. In other words, we assume that at most one fork occurs per round. By modeling the blockchain network based on our concept of rounds, we can more accurately account for the impact of blockchain forks, thereby enabling a much more precise calculation of mining fairness.
@@
Subsequently, we validate the accuracy of our proposed model-based calculation in measuring mining fairness by conducting simulation experiments. However, validating our method in large-scale networks would be challenging due to computational constraints. Hence, we perform the validation in networks with small numbers of miners. Our results demonstrate that the model-based calculation quantitatively determines mining fairness much more accurately than existing methods.
In Section 2, we introduce the concept of rounds. It becomes possible to achieve a more accurate calculation of mining fairness by capturing the impact of forks on rounds. Section 3 discusses related work. Section 4 describes the proposed method, the model-based calculation. Section 5 presents the validation of the proposed method. Section 6 provides the conclusion.
@@ 2 Rounds
Herein, we introduce the concept of rounds in a blockchain to accurately capture and incorporate the impact of forks. A round is a time interval defined from a global perspective. Specifically, round $r$ refers to the time from the first generation of a block at height $r$ to the first generation of a block at height $r+1$. The height of round $r$ is defined as $r$.
@@
Because forks occur probabilistically within a blockchain network, the number of blocks per round does not always equal one. After the first block at height $r$ has been generated, another miner may generate a new block before the first block is propagated throughout the network. Notably, not all blocks generated in round $r$ have a height of $r$. For example, a miner unaware of the block at height r1r-1 will generate a block at height r1r-1 during round $r$ (Fig. 2).
@@
The round start rate of each miner is defined as the probability that the miner initiates a round. Note that, owing to the occurrence of forks, the round start rate does not equal the proportion of the hashrate.
@@
With the introduction of rounds, we can formally define the fork rate. The fork rate is the probability that the number of blocks per round is two or more. When the number of blocks per round is two or more, the blockchain diverges or forks. Therefore, forks (and fork rates) are named as such because they function similarly to a traditional fork.
3 Related Work
3.1 Mining Fairness for Unintentional Forks
Croman et al. proposed several metrics and solutions to address the scalability problems of blockchains [10]. They highlighted mining fairness as a potentially more valuable metric despite its difficulty in measurement.
Since then, numerous studies have attempted to analyze mining fairness [11, 12, 13, 14, 15, 16, 17]; however, none have accurately captured the impact of forks on mining fairness. Consequently, the analyses conducted thus far diverged from the actual values observed in blockchain networks. Herein, we discuss each of these studies in detail.
Kanda et al. introduced the concept of effective hashrate, based on the idea that miners cannot contribute to the main chain until they receive the latest block [11]. They calculated the effective hashrate by multiplying the original hashrate by the ratio of the time required to receive a block to the block generation interval. They asserted that the proportion of the effective hashrate equals the block reward rate. Jiang et al. analyzed mining fairness based on a concept similar to that proposed by Kanda et al. [12]. They calculated the average block reception time for each miner and defined the maximum difference in the average block reception times as mining fairness. However, their concept of mining fairness evidently deviates from reality.
Xiao et al. proposed a model-based approach to analyzing mining fairness in a blockchain network to analyze the blockchain network connectivity [14]. The main difference between their approach and our model-based calculation is that the round start rate is considered in the proposed method. Conversely, they assumed that the round start rate equals the proportion of hashrate, which is not necessarily true.
Mao et al. investigated how the manner in which miners are connected affects mining fairness [17]. They conducted both theoretical analyses and simulations; however, each approach had limitations. Regarding their theoretical analysis—they considered only a few blocks directly connected to the genesis block. Conversely, regarding their simulation analysis—they ran a scenario where the ratios of the propagation time to the average block generation interval, $d/T$, were extremely high. Although this condition might allow the observation of trends related to mining fairness, it does not accurately reflect the behavior of actual blockchain networks, raising concerns about the generalizability of the results.
Huang et al. compared the proof of work (PoW) and proof of stake (PoS) from the perspective of mining fairness [13]. They claimed that mining fairness is achieved in PoW blockchains in that the block reward rate equals the proportion of the hashrate. However, this conclusion arose because they did not consider unintentional forks in PoW blockchains. They also analyzed the convergence rate in addition to the expected value of mining fairness. Conversely, in this study, we analyzed only the expected value of mining fairness.
3.2 Mining Fairness for Intentional Forks
Attacks that intentionally compromise mining fairness to increase block reward rates unjustly have been studied extensively [18, 5, 19, 20]. For example, Eyal et al. proposed a mining strategy known as selfish mining, which increases the block reward rate by intentionally causing forks [5]. They defined the success of selfish mining as achieving a block reward rate that exceeds the proportion of the hashrate, indicating that local mining fairness becomes positive. Sun et al. introduced the Kullback–Leibler (KL) divergence between the distributions of block reward rates and proportions of hashrate as a measure of the impact of selfish mining [21]. This concept is similar to global mining fairness. However, while KL divergence measures the difference between distributions, we are interested in the distribution of the difference between block reward rates and the proportions of hashrate. The KL divergence can be zero even when mining fairness is compromised. Therefore, their definition of mining fairness has limited expressiveness. Consequently, this study does not address mining fairness based on how they defined it.
@@ 4 Model-Based Calculation
We propose a model-based calculation of mining fairness. In this approach, we replace the complex real-world blockchain network with a simpler model in which at most two blocks can be generated per round. Here, a “round” refers to the concept introduced in Section 2. This simplification allows us to appropriately account for the impact of forks on mining fairness, thereby significantly improving the accuracy of mining fairness calculation.
@@ 4.2 Definition of Mining Fairness
Before calculating mining fairness, we first define it. In this study, mining fairness is divided into local mining fairness and global mining fairness. We define local mining fairness, $LF$, based on two measures, as follows:
@@
Local mining fairness is particularly useful for individual miners, while global mining fairness is important for system designers and engineers. For instance, miners aim to select the most profitable strategies, which is inherently equivalent to improving local mining fairness. On the other hand, system designers seek to establish a fair mining ecosystem, making global mining fairness a crucial objective.
@@ 4.3 Calculation of Mining Fairness
Let $X_{r}$ be a random variable representing the miner that generates the block that starts the round $r$. Then, the following equation holds:
$P(X_{r+1}=i)=$
$\sum_{j\in V}\left(\alpha_{i}(1-F_{ji})+\sum_{k\in V}\alpha_{k}F_{jk}\alpha_{i}\right)P(X_{r}=j).$ (5)
Notably, $P(X_{r+1}=i)$ is only dependent on $P(X_{r}=j)$. Therefore, the stochastic process $\{X_{r}\}_{r=0}^{\infty}$ is a Markov chain. Additionally, this Markov chain is ergodic in most cases because $F_{ij}$ is less than 1 and $\alpha_{i}(1-F_{ji})+\sum_{k}\alpha_{k}F_{jk}\alpha_{i}$ is usually positive. Consequently, a unique stationary distribution exists, and the limit distribution is stationary. We can then determine the stationary distribution by iterating (5).
Let the limit distribution be $\pi$. This represents the distribution of the miners that generate blocks that start rounds after sufficient time has passed. Using $\pi$, the block reward rate $r_{i}$ for each miner is given by the following equation:
@@ 4.4 Algorithm
In this section, we describe the algorithm used in this paper study to calculate mining fairness, as detailed in Section 4.3.
Algorithm 1 employs an iterative method to compute the mining fairness for each miner. The round start rate calculation is performed between lines 9 and 27. Specifically, the fork rate is precomputed between lines 10 and 16. The variable loop manages the operations executed in each iteration. The calculations within the for loop from lines 20 to 26 follow the same process as described in (5). Mining fairness is computed between lines 28 and 34, with the for loop calculations corresponding to Equations (8) and (9).
@@
2:$V$ $\triangleright$ set of miners
3:NN $\triangleright$ number of miners
[…]
5:$F[N][N]$ $\triangleright$ fork rate
6:$W[N][N]$ $\triangleright$ winning rate
@@
[…]
@@
[…]
15:$dp[N]$ $\triangleright$ for dynamic programming
@@
23:while $\exists i\in V$ s.t. $|\pi[i][loop\ \bmod\ 2]-\pi[i][(loop+1)\ \bmod\ 2]|>\epsilon$ do
24:  $loop\leftarrow(loop+1)\ \bmod\ 2$
@@
28:     $\pi[i][(loop+1)\ \bmod\ 2]\leftarrow\pi[i][(loop+1)\ \bmod\ 2]+\alpha[i](1-F[j][i])\pi[j][loop]$
29:     $\pi[i][(loop+1)\ \bmod\ 2]\leftarrow\pi[i][(loop+1)\ \bmod\ 2]+dp[j]\alpha[i]\pi[j][loop]$
@@
38:   $LF_{1}[i]\leftarrow LF_{1}[i]+\pi[j]\alpha[i]F[j][i](1-W[j][i])-\pi[i]\alpha[j]F[i][j](1-W[i][j])$
@@ 4.5.1 How to Determine $F_{ij}$
4.5.2 Tips for How to Determine $W_{ij}$
Prior to any explanations regarding $W_{ij}$, first discussing the concept of chain ties is crucial.
Each miner constructs chains from their blocks and selects one main chain among them. The rule for selecting this chain is known as the fork choice rule. For instance, in Bitcoin, the longest chain rule, which selects the longest chain, is adopted.
However, in some cases, the fork choice rule alone may not uniquely determine the main chain owing to the occurrence of forks. This situation is known as a chain tie. A tie-breaking rule is implemented to resolve a chain tie. For instance, in Bitcoin, the first-seen rule, which selects the chain received first, is adopted.
@@ 4.5.2 Tips for How to Determine $W_{ij}$
Selects the latest chain among the chains in a tie [22, 23, 24]. Suppresses selfish mining more effectively than the random rule.
@@ 5 Validation
Herein, we validate the capability of our proposed model-based calculation to determine mining fairness accurately. First, we examine the assumption that the number of blocks per round is at most two from the perspective of the scale of the fork. Next, we compare the results of simulation experiments with those of the model-based calculation.
While it is excessively time-consuming to calculate mining fairness via simulations of networks composed of many (approximately 100 or more) miners, it is feasible to calculate mining fairness accurately and relatively quickly for networks with a small number of miners (2–10). In this study, we perform the validation using networks comprising two and ten miners, demonstrating that the proposed model-based calculation determines mining fairness much more accurately than existing methods.